OpenStudy (anonymous):

find the magnitude and direction angle of v = -4i - 2j

OpenStudy (tkhunny):

This should ring some bells. \(|v| = \sqrt{(-4)^{2}+(-2)^{2}}\) What about that direction?

OpenStudy (anonymous):

i is the x direction, and j is the y direction.

OpenStudy (anonymous):

If that was to be plotted on a standard x-y axis plane.

OpenStudy (anonymous):

i solved the equation up above and I got sqrt of 20..don't know where to go from here

OpenStudy (tkhunny):

Ever hear of a "Reference Angle"?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

What you solved above, that will give you the magnitude.

OpenStudy (anonymous):

2 sqrt of 5 is the same as the sqrt of 20, yes?

OpenStudy (tkhunny):

Excellent. Thinking on a Right Triangle, pick an acute angle. Once you have done that, label the adjacent leg "4" and the opposite leg "2". Now what?

OpenStudy (tkhunny):

Yes, \(\sqrt{20} = 2\sqrt{5}\)

OpenStudy (anonymous):

now i have to use the Pythagorean theorem to find the missing side?

OpenStudy (tkhunny):

If you would like, but it's quicker just to use the inverse tangent to find the reference angle you need.

OpenStudy (anonymous):

I got 207 as the direction

OpenStudy (tkhunny):

Keep in mind that in a Right Triangle, we will get ONLY a reference angle. It's your responsibility to add 180º to get it into Quadrant III where it belongs.

OpenStudy (tkhunny):

Right, 206.565º